one mole of n2 and 3 moles of h2 are placed in a closed container at a pressure of 4 atm. The pressure falls to 3 atm at the same temperature when N2 + 3H2 gives 2NH3. the equilibrium constant for dissociation of NH3 is

Grade:11

2 Answers

Vidyashree
6 years ago
                   N2 + 3H2   ---------> 2NH3
Initial            1        3                      0              P1=4 atm   ∑n = 4
Eqbm         1-x     3-3x                  2x             P2=3 atm   ∑n = 4-2x
 
P α n
P1/P2 = n1/n2
4/3 = 4/(4-2x)
x= 0.5
 
PN2 = 0.5,   PH2 = 1.5,  PNH3 = 1
 
K = ( PNH3)2 / (PN2) (PH2)^3
K= 1 / (0.5) (1.5)^3
the equilibrium constant for dissociation of NH3 is K = (0.5) (1.5)^3
 
 
ankit singh
3 years ago
 N2 + 3H2   ---------> 2NH3
Initial            1        3                      0              P1=4 atm   ∑n = 4
Eqbm         1-x     3-3x                  2x             P2=3 atm   ∑n = 4-2x
 
P α n
P1/P2 = n1/n2
4/3 = 4/(4-2x)
x= 0.5
 
PN2 = 0.5,   PH2 = 1.5,  PNH3 = 1
 
K = ( PNH3)2 / (PN2) (PH2)^3
K= 1 / (0.5) (1.5)^3